An Airbus A380-800 passanger airplane is cruising at constant altitude on a straight line with a constant speed. The total surface area of the two wings is 395 m^2. The average speed of the air just below the wings is 248 m/s, and it is 276 m/s just above the surface of the wings. What is the mass of the airplane? The average density of the air around the airplane is ρair = 1.29 kg/m^3.
Answer #1

Using Bernoulli's Principle:

Pu + 0.5*rho*Vu^2 = Pl + 0.5*rho*Vl^2

Pl - Pu = 0.5*rho*(Vu^2 - Vl^2)

Also,

Pl - Pu = F/A = mg/A

mg/A = 0.5*rho*(Vu^2 - Vl^2)

m = A*0.5*rho*(Vu^2 - Vl^2)/g

Using given Values:

m = 395*0.5*1.29*(276^2 - 248^2)/9.81

m = 381045.75 Kg